OpenStudy (anonymous):

Solve for x. 5x - 11 = 7x - 3

OpenStudy (uri):

5x - 11 = 7x - 3 Add 11 both sides 5x - 11 = 7x - 3 +11 +11 =========== 5x=7x+8

OpenStudy (nurali):

5x - 11 = 7x - 3 -11+3=7x-5x -8=2x x=-4

OpenStudy (uri):

5x-7x=8 -2x=8 x=8/-2 x=-4 ans